What volume of Cl2(g) measured at STP is produced when 7.65 g of HCl(aq) reacts?

1
Expert's answer
2022-05-03T02:04:04-0400

H2 + Cl2 = 2HCl

V(Cl2) = ( m(HCL) * Vm) / 2Mr(HCl)

V(Cl2) =( 7.65 g * 22.4 L/mol) / 2 * 36.5 g/mol = 2.35 L